approximate the (a) smaller and (b) larger wavelength at which the eye of a standard observer has half the eye’s maximum sensitivity. What are the (c) wavelength, (d) frequency, and (e) period of the light at which the eye is the most sensitive?
Answer #1

(a) Smaller wavlenght is 515 nm

(b) larger wvaelenght is 610 nm

(c) the wavelength which eye is most sensitive is 555 nm

(d) f =c/lamda = 3x108/ 555x109

f = 5.41x1014 Hz

(e) T =1/f = 1/5.41x1014

T= 1.85x10-15 s